OpenStudy (anonymous):

Is this proof that 1+1=3?

OpenStudy (anonymous):

where is proof!lol

OpenStudy (anonymous):

1+1=3??????????????????

OpenStudy (anonymous):

A friend agave me this puzzling proof that 1=1=3. It sounds stupid but look: x = y -x2= -xy x2 - y2 = x2 - xy (x + y)(x - y) = x(x - y) x + y = x 1 + 1 =2 2 = 1 I can't believe it! PS the "2" at the end is a "squared."

Parth (parthkohli):

Well, the thing here is actually, when your friend goes from Step 4 to 5, he divides by zero.

Parth (parthkohli):

I have seen a lot of proofs like this. All of them are wrong.

OpenStudy (anonymous):

OK ... BOss

Parth (parthkohli):

Dividing by zero ==> INFINITE

OpenStudy (anonymous):

Well, I disagree that it's infinite. It's more accurate to say that it's undefined.

Parth (parthkohli):

Both apply, in a way.

OpenStudy (anonymous):

I mean, conceptually it can be helpful to think of it as infinity, but in reality, it's more accurate to say undefined. Here's my example. We know that Distance = rate * time. Therefore, if I know how far a trip is and how fast I'm traveling, I can determine how long it takes to get there. Distance/rate = time So, suppose the trip is 10 miles, and I will travel 0 mph. How long will it take me to get there? 10/0... You could think of it as "infinity," or it's taking you forever to get there, but it's more accurate to say "Undefined," which can be understood as "there's no answer" or "at no time." In other words, you WON'T get there.